Five unbiased dice were thrown 96 times and the number of times 4, 5 or 6 was 

obtained, is given in the following table:

 

No. of dice  0. 1. 2. 3. 4. 5.

showing 4, 

5 or 6 


Frequency 1 10 24 35 18 8


At 5% level of significance test whether this data comes from a binomial distribution. 

You may like to use the following values. 

 [( x5210.05)=11.07 , x62(0.05)=12.59 ,x72(0.05)=14.07.](

1
Expert's answer
2021-06-09T10:32:03-0400

Let H"_o": The given data comes from Binomial distribution

Total number of dice is thrown N=96 times

Probability of obtaining 4,5 or 6 "p=\dfrac{1}{2}"

"q=1-p=1-\dfrac{1}{2}=\dfrac{1}{2}"


For binomial distribution:

"E_i=Np_i=N\/f_i"

where pi is the probability of success,

fi is the frequency of success.


The table for Chi-square is given by,-




The value of "\chi^2=\sum \dfrac{(O_i-E_i)^2}{E_i}=14.01"


The tabulate value of "\chi^2" at 5% level of significance is 14.07.


Conclusion: As the calculated value of "\chi^2" is less than the tabulated value So H_o is accepted i.e Data comes from binomial distribution.